Let f left(x right)=x^2−3x+3
find f(-1)
a.-7
b.7
c.4
d.1
1
Expert's answer
2020-05-15T17:02:16-0400

"f(-1)=(-1)^2-3(-1)+3=1+3+3=7."

Var. b.